# Concerns and Values in Human-Robot Interactions: A Focus on Social Robotics

### v1

@@ Abstract
Robots, as AI with physical instantiation, inhabit our social and physical world, where their actions have both social and physical consequences, posing challenges for researchers when designing social robots. This study starts with a scoping review to identify discussions and potential concerns arising from interactions with robotic systems. Two focus groups of technology ethics experts then validated a comprehensive list of key topics and values in human-robot interaction (HRI) literature. These insights were integrated into the HRI Value Compass web tool, to help HRI researchers identify ethical values in robot design. The tool was evaluated in a pilot study. This work benefits the HRI community by highlighting key concerns in human-robot interactions and providing an instrument to help researchers design robots that align with human values, ensuring future robotic systems adhere to these values in social applications.
@@ 1 Introduction
Recent advancements in artificial intelligence and robotics are pushing robots from research laboratories into public spaces [61], healthcare facilities [52] and schools [12], as well as private homes. As robots start exhibiting nuanced, multifaceted behaviours, by inhabiting high-stakes environments where their actions have profound social and physical consequences, robots become more and more social agents [9, 3]. As such, they transcend their nature of complex tools and start to pose new ethical challenges, impacting the humans involved in immediate and long-term interactions.
@@
Human-robot interaction (HRI) researchers must anticipate and address these ethical concerns. This involves understanding the potential risks and benefits associated with robot interactions and striving to embed ethical principles and values into the design of robotic systems that align with societal norms and values. In social psychology, seminal works such as Schwartz’s theory of basic human values [82] have pursued a categorisation of values. Similarly, HCI literature demonstrates how ethical frameworks, like Design Justice [24], can guide the creation of ethical technologies.
@@
The HRI field has investigated the ethical challenges involved in its specific context with existing frameworks [67]. These tools elicit the values of stakeholders – such as roboticists, healthcare workers, and teachers – and embed them into the design of robotic systems. However, the emergent ethical concerns arising from the interactions themselves and the relationships between humans and robots in the short and long term are still underexplored.
@@
These interaction-driven values are ethical considerations that emerge during human-robot interactions, such as fairness, dependency, and autonomy. They are context-sensitive and evolve over time, making them distinct from predefined stakeholder values. Addressing them is crucial as they shape user experiences and the societal impact of robotic systems. Indeed, only recently the technology has matured enough to allow overcoming challenges such as complex real-world dynamics. Furthermore, practical tools are still needed to help researchers identify and integrate these interaction-driven values into robotic technologies.
@@
Aiming to fill the gap in the HRI literature,we investigate RQ1: what discussions on values have emerged in HRI literature to date? We scope our search to applications of robots for social good, in the context of healthcare, education, and private homes. This limitation is justified by the significant societal impact and ethical complexity inherent in these contexts, where robots are increasingly integrated into personal and sensitive domains, directly influencing human wellbeing and quality of life. In addition, we focus on the values arising from the interactions between humans and robots, rather than the stakeholders’ values, as the latter has already been widely explored [108, 67, 87].
@@
The HRI field is inherently multidisciplinary. This can make addressing the diverse needs and perspectives from various areas of expertise challenging. Recently, philosophers and ethics experts have become more involved in the robotic design process [7]. However, the HRI literature has yet to fully explore the potential benefits of involving these stakeholders in defining HRI values. Therefore, our second research question is RQ2: what values can we define in HRI with the help of experts in the field?
@@
Lastly, HRI researchers often struggle to incorporate the complex and multifaceted aspects of ethics into their robot design interactions due to the complexity of the field, lack of expertise in this area [34], and absence of adequate tools that help researchers include interaction-driven ethical aspects and values in their design process [54]. Hence, our third research question is RQ3: how can we assist HRI researchers in identifying values in design scenarios?
@@
To address these questions, our work presents a closed-loop process consisting of three main phases as depicted in Figure 1. In Phase 1 – Section 3 – we conducted a scoping review to identify the values discussed in HRI literature over the last two decades, starting from the HRI field’s emergence in 2006 [2]. In Phase 2 – Section 4 – we organised two focus groups with experts in philosophy and ethics of technology to review these values and establish a set of topics and values relevant to HRI contexts, reported in Section 4.3. Finally, in Phase 3 – Section 5 – we carried out a user study to evaluate a tool [^1] Available online at https://hri-value-compass.github.io/ we developed from the outcomes of the previous phases, the HRI Value Compass, designed to help HRI researchers identify relevant values when designing robotic interfaces. The tool references the relevant literature, offering additional insights to the researchers, and closing the process loop.
@@ 2.1 Values
Schwartz [81] defines values as deeply held beliefs that are closely tied to emotions and serve to motivate actions toward desirable goals, such as independence or justice. Unlike norms or attitudes, values transcend specific actions or situations and act as guiding standards for evaluating behaviour, policies, or events. Although people may not always be consciously aware of their values, these values influence decisions, particularly when they come into conflict with one another.
Schwartz’s theory organises values into four broad categories: openness to change, self-enhancement, conservation, and self-transcendence. Each of these higher-order categories includes specific values, such as self-direction, power, tradition, and universalism, resulting in a total of 19 values in the revised model [82]. Although Schwartz’s framework is valuable, it may not fully capture the context-specific values that arise in HRI, where interactions between humans and robots introduce unique ethical considerations.
Graham et al. [41] propose a model of moral foundations, which are psychological systems underlying human moral reasoning. These foundations are divided into six categories: care/harm, fairness/cheating, loyalty/betrayal, authority/subversion, sanctity/degradation, and liberty/oppression. While this model provides insights into moral reasoning, its high-level abstractions limit its direct application in the design of technological systems, including those in HRI. As such, identifying values specifically relevant to HRI, which may not be fully encompassed by these general models, remains crucial.
Another definition of values is the one presented by Scanlon [80] in the buck-passing theory. In his discourse about values, Scanlon states that “to call something good is to claim that it has other properties […] which provide such reasons” for behaving in certain ways with respect to it. The value is thus not an intrinsic property of an object or situation, but entirely depends on “such reasons”. We argue that since this definition imposes less constraints, it is better suited to model the nuanced dynamics taking place in HRI. Consequently, the definition of value that we adopt does not follow a universal maxim and instead is more attuned to contractualist theories [79], which emphasise respect and consideration of individual preferences in ethical decision-making.
@@ 2.2 Stakeholder perspectives
Several studies have explored the perspectives of stakeholders in high-stakes environments, providing valuable insights into the ethical challenges of HRI. Cheon and Su [22] highlight the importance of including roboticists’ perspectives in their design, particularly when addressing ethical concerns. In support of this, toolkits and frameworks have been developed to assist researchers in identifying and integrating ethical values into their design processes [77]. Design methodologies, such as Design Justice [24], seek to address biases in technology design and ensure that systems are developed in a manner that is fair and just.
@@
In the field of HCI, more concrete ethical concerns have been identified. For example, Makasi et al. [56] examined chatbots providing public services and identified 14 key public service values. Although some of these values are specific to the public service context, many – such as adaptability, fairness, accountability, and privacy – are more broadly applicable to AI systems, including AI-powered robots. These ethical considerations offer important insights for the design of systems that interact directly with humans, such as social robots.
In the context of robotics, Fosch-Villaronga et al. [34] analysed discussions with experts, revealing key ethical, legal, and social challenges in the deployment of social robots. The authors identified five primary areas of concern: security and privacy, legal uncertainty, autonomy and agency, employment, and the replacement of human interaction. These concerns are particularly pertinent to robots designed for social good, such as those used in healthcare, education, and domestic settings. In our work, we aim to build on these insights by identifying concerns and values specific to HRI, providing a comprehensive list of values that can guide the design of robotic systems that align with human values and societal norms.
Ostrowski et al. [67] conducted a comprehensive review of HRI literature over the past two decades, examining various interpretations of ethics within the field. One particularly relevant aspect of their work is the exploration of ethics in human-robot relationships. In this context, they successfully identified several key ethical themes, which we aim to further validate and expand upon in our own research.
@@ 3 Phase 1: Scoping Review
The main objective of this phase is to understand the key discussions revolving around values that have emerged in the HRI field in the last two decades (RQ1). To achieve this, we conducted a scoping review focused on identifying and extracting the concerns and values from HRI studies published since 2006, when the ACM/IEEE International Conference on Human-Robot Interaction was established [2].
@@ 3.1 PRISMA process
We followed the PRISMA framework [68], the state-of-the-art workflow for systematic literature reviews, adapted for a scoping review [95]. The process includes the following steps, as shown in Figure 2: (i) identification, searching specific databases within a defined time range, (ii) screening, filtering out records whose titles and abstracts do not meet the eligibility criteria, (iii) selection, reading the full texts of the publications and excluding those that do not meet the eligibility criteria, and (iv) inclusion, selecting the final list of publications and extracting their data.
@@ 3.1.1 Search query
Based on the research question RQ1 and the scope of this work, we constructed a search query as in [21, 91]. The query is satisfied for a record if the combination of the record’s title, abstract and keywords satisfies the following: contains the term social; contains the term robot or its variations; contains any of the terms value, ethic, moral, deont- (for deontology); and contains any of the terms school, teach, education, assist, healthcare, patient, home, house, domestic.
@@ 3.1.2 Exclusion criteria
  • no physical robot was present (e.g.: virtual agents are excluded);
@@ 3.1.3 Selection process
Subsequently, the retrieved publications were again split randomly into two sets and assigned to the same reviewers, who read the full texts for the final selection and data extraction. Again, if a reviewer was uncertain about including a paper, this was discussed with the other reviewer for a final decision. The final selection comprises 43 publications and a list of the publications is available in Tables 2 and 3 and also in a publicly accessible table stored in our web tool [^2] https://hri-value-compass.github.io/.
@@ 3.1.4 Data extraction and analysis
Variable Type Example
Publication’s year Numerical 2010, …
Application topic Categorical healthcare, education, …
Authors’ affiliation country Categorical USA, Germany, Europe, …
Empirical study Boolean 0, 1
Population’s country Categorical United Kingdom, …
Population’s age Categorical Children, Elderly, …
Transcripts Qualitative “the issues of infantilisation of elderly […]”
@@
For extracting the qualitative variables, we adopted the thematic analysis approach [13], which consists of the following six steps. During the data extraction phase described above, for each paper assigned to them, each of the reviewers (Step 1) familiarised themselves with the data by copying verbatim parts of the papers (transcripts) that were relevant to the scope. Subsequently, they (Step 2) generated the initial codes and (Step 3) searched for themes by collating codes. Then, together, they (Step 4) reviewed the themes, comparing their results and discussing the differences, (Step 5) named the themes into meaningful values according to the literature, and (Step 6) produced a report.
@@ 3.2 Results: value-related HRI papers
Table 2: List of papers included in our survey and corresponding values and topics extracted. Note that this table has been updated with the final values defined at the end of the design process to avoid any confusion. The initial list of values can be found in the Appendix D [Continued in Table 3].
@@
Authors Year Application Study Topics Values Jung et al. [51] 2023 manipulation yes companionship, exclusion, education, manipulation, connectedness, equity, welfare, transparency, attachment autonomy, privacy Wagman and Parks [103] 2021 feminism no exclusion, bias, manipulation, fairness, expectations equity, autonomy Van Ewijk et al. [97] 2020 education yes wellbeing, accountability, manipulation, transparency, welfare, equity, autonomy, consent, social substitute privacy, connectedness Li et al. [55] 2019 home yes dilemma, meta N/A Huber et al. [46] 2016 attachment no profiling, dependence, identity disclosure privacy, agency, dignity Ostrowski et al. [66] 2022 healthcare yes independence, attachment, profiling, transparency agency, autonomy, privacy Tuuli and Jaana [96] 2020 healthcare yes meta, expectations N/A Malle et al. [58] 2020 ethics no meta N/A Poulsen et al. [69] 2018 trust yes identity disclosure, transparency dignity, privacy Amirabdollahian et al. [6] 2013 healthcare no independence, safety, social enabler agency, welfare, connectedness Yeung et al. [109] 2023 ethics yes social substitute, wellbeing, nonjudgmental, mentor, connectedness, welfare, virtue social enabler Rosero [75] 2023 trust yes dilemma, identity disclosure dignity Soares et al. [88] 2023 healthcare yes independence, accountability, expectations agency, equity Felber et al. [31] 2022 healthcare no exclusion, profiling, accountability, equity, privacy, dignity, degradation, social substitute connectedness Carros et al. [20] 2022 ethics yes dilemma, meta, expectations N/A Street et al. [93] 2022 healthcare yes independence, profiling, identity disclosure, agency, privacy, dignity, dependence, corruptor, control, social substitute virtue, connectedness Malinverni et al. [57] 2021 education yes wellbeing welfare Smakman et al. [87] 2021 education yes education, exclusion, independence, profiling, wellbeing, welfare, equity, agency, accountability, access, manipulation, attachment, privacy, autonomy, virtue, transparency, fairness, mentor, connectedness expectations, consent, social substitute Vanderelst and Willems [100] 2020 healthcare yes accountability, expectations, profiling, control equity, privacy, agency Vandemeulebroucke et al. [99] 2020 healthcare no exclusion, independence, profiling, social enabler, agency, equity, privacy, wellbeing, accountability, bias, dependence, intimacy, connectedness, welfare, dignity tricking, degradation, control, safety, meta, social substitute Cappuccio et al. [19] 2021 moral patiency no exclusion, bias, dependence, corruptor, equity, agency, virtue fairness, mentor, expectations Poulsen et al. [70] 2020 healthcare yes safety, profiling privacy, welfare Battistuzzi et al. [10] 2018 healthcare no tricking, intimacy, independence, companionship dignity, privacy, agency, connectedness Čaić et al. [17] 2018 healthcare yes companionship, profiling, social enabler, connectedness, privacy, agency, dependence, assistance, safety, social substitute welfare Fiske et al. [33] 2019 mental no profiling, accountability, tricking, manipulation, privacy, equity, dignity, attachment, meta, expectations, social substitute autonomy, connectedness Cresswell et al. [25] 2018 healthcare yes tricking, safety, control dignity, welfare, agency Draper and Sorell [30] 2017 healthcare yes wellbeing, accountability, dependence, welfare, equity, agency, access, manipulation, control, meta, privacy, autonomy, connectedness expectations, social substitute Ienca et al. [48] 2016 healthcare no safety, meta, transparency welfare, privacy Sharkey [85] 2016 healthcare no profiling, identity disclosure, attachment, privacy, dignity, autonomy, transparency, social substitute connectedness
@@
Table 3: [Continued from Table 2] List of papers included in our survey and corresponding values and topics extracted. Note that this table has been updated with the final values defined at the end of the design process to avoid any confusion. The initial list of values can be found in the Appendix D.
@@
Authors Year Application Study Topics Values García-Soler et al. [37] 2018 healthcare yes companionship connectedness Amirabdollahian et al. [5] 2013 safety, meta, profiling welfare, privacy Calo et al. [18] 2011 healthcare yes social substitute, tricking, manipulation, connectedness, dignity, autonomy attachment, meta, expectations, social enabler Van Wynsberghe [98] 2013 ethics no accountability, meta, control equity, agency Fridin [36] 2014 education yes safety, profiling, attachment welfare, privacy, autonomy Sharkey and Sharkey [84] 2012 healthcare no independence, intimacy, tricking, safety, social substitute agency, privacy, dignity, connectedness, welfare Sharkey and Sharkey [83] 2010 healthcare no social substitute, intimacy, tricking, connectedness, privacy, dignity, control, social enabler agency Misselhorn et al. [62] 2013 healthcare no tricking, degradation, wellbeing, social substitute dignity, welfare, connectedness Hung et al. [47] 2022 healthcare yes consent, social enabler connectedness, privacy Baisch and Kolling [8] 2022 healthcare yes bias, expectations, attachment equity, autonomy Wheeler [105] 2022 healthcare no exclusion, wellbeing, transparency welfare, privacy Prescott and Robillard [71] 2021 relationship no companionship, social substitute, wellbeing, connectedness, welfare , dignity degradation, safety, expectations, social enabler Smakman et al. [86] 2020 education yes education, profiling, accountability, tricking, welfare, privacy, equity, manipulation, control, nonjudgmental, attachment, dignity, autonomy, agency, assistance, mentor, social substitute virtue, connectedness Aldinhas Ferreira and Sequeira [4] 2017 healthcare no intimacy, profiling, social substitute privacy, connectedness

### v2

@@ Abstract
Robots, as AI with physical instantiation, inhabit our social and physical world, where their actions have both social and physical consequences, posing challenges for researchers when designing social robots. This study starts with a scoping review to identify discussions and potential concerns arising from interactions with robotic systems in the context of healthcare, education, and private homes. Two focus groups of technology ethics experts then validated a comprehensive list of key topics and values in human-robot interaction (HRI) literature in these contexts. These insights were integrated into the HRI Value Compass web tool, to help HRI researchers identify these values in robot design. The tool was evaluated in a pilot study. This work benefits the HRI community by highlighting key concerns in human-robot interactions and providing an instrument to help researchers design robots that align with human values, ensuring future robotic systems adhere to these values in social applications.
@@ 1 Introduction
Recent advancements in artificial intelligence and robotics are pushing robots from research laboratories into public spaces [mintrom2022robots], healthcare facilities [kyrarini2021survey] and schools [benitti2012exploring], as well as private homes. As robots start exhibiting nuanced, multifaceted behaviours, by inhabiting high-stakes environments where their actions have profound social and physical consequences, robots become more and more social agents [bartneck2024human]. As such, they transcend their nature of complex tools and start to pose new ethical challenges, impacting the humans involved in immediate and long-term interactions.
@@
Human-robot interaction (HRI) researchers must anticipate and address these ethical concerns. This involves understanding the potential risks and benefits associated with robot interactions and striving to embed ethical principles and values into the design of robotic systems that align with societal norms and values. In social psychology, seminal works such as Schwartz’s theory of basic human values [schwartz_refining_2012] have pursued a categorisation of values. Similarly, HCI literature demonstrates how ethical frameworks, like Design Justice [costanza2020design], can guide the creation of ethical technologies.
@@
The HRI field has investigated the ethical challenges involved in its specific context with existing frameworks [ostrowski2022ethics]. These tools elicit the values of stakeholders – such as roboticists, healthcare workers, and teachers – and embed them into the design of robotic systems. However, the emergent ethical concerns arising from the interactions themselves and the relationships between humans and robots in the short and long term are still underexplored.
@@
These interaction-driven values are ethical considerations that emerge during human-robot interactions, such as fairness, dependency, and autonomy. They are context-sensitive and evolve over time, making them distinct from predefined stakeholder values. Addressing them is crucial as they shape user experiences and the societal impact of robotic systems. Indeed, only recently has the technology matured enough to allow overcoming challenges such as complex real-world dynamics. Furthermore, practical tools are still needed to help researchers identify and integrate these interaction-driven values into robotic technologies.
@@
Aiming to fill the gap in the HRI literature, we investigate RQ1: what discussions on values have emerged in HRI literature to date? We scope our search to applications of robots for social good, in the context of healthcare, education, and private homes. This limitation is justified by the significant societal impact and ethical complexity inherent in these contexts, where robots are increasingly integrated into personal and sensitive domains, directly influencing human wellbeing and quality of life. In addition, while keeping into consideration the previous literature on human values, we adopt as a starting point the values arising from the interactions between humans and robots, rather than the stakeholders’ values, as the latter has already been widely explored [wullenkord2020societal, ostrowski2022ethics, smakman2021moral].
@@
The HRI field is inherently multidisciplinary. This can make addressing the diverse needs and perspectives from various areas of expertise challenging. Recently, philosophers and ethics experts have become more involved in the robotic design process [axelsson2024robots]. However, the HRI literature has yet to fully explore the potential benefits of involving these stakeholders in defining HRI values. Therefore, our second research question is RQ2: what values can we define in HRI with the help of experts in the field?
@@
Lastly, HRI researchers often struggle to incorporate the complex and multifaceted aspects of ethics into their robot design interactions due to the complexity of the field, lack of expertise in this area [fosch2020gathering], and absence of adequate tools that help researchers include interaction-driven ethical aspects and values in their design process [leveringhaus2018developing]. Hence, our third research question is RQ3: how can we assist HRI researchers in identifying values in design scenarios?
@@
To address these questions, our work presents a closed-loop process consisting of three main phases as depicted in Figure 1. In Phase 1 – Section 3 – we conducted a scoping review to identify the values discussed in HRI literature over the last two decades, starting from the HRI field’s emergence in 2006 [10.1145/1121241], with a focus on healthcare, education, and private homes. In Phase 2 – Section 4 – we organised two focus groups with experts in philosophy and ethics of technology to review these values and establish a set of topics and values relevant to these HRI contexts, reported in Section 4.3. Finally, in Phase 3 – Section 5 – we carried out a user study to evaluate a tool [^1] Available online at https://hri-value-compass.github.io/ we developed from the outcomes of the previous phases, the HRI Value Compass, designed to help HRI researchers identify relevant values when designing robotic interfaces. The tool references the relevant literature, offering additional insights to the researchers, and closing the process loop.
@@ 2.1 Values
schwartz_overview_2012 defines values as deeply held beliefs that are closely tied to emotions and serve to motivate actions toward desirable goals, such as independence or justice. Unlike norms or attitudes, values transcend specific actions or situations and act as guiding standards for evaluating behaviour, policies, or events. Although people may not always be consciously aware of their values, these values influence decisions, particularly when they come into conflict with one another.
Schwartz’s theory organises values into four broad categories: openness to change, self-enhancement, conservation, and self-transcendence. Each of these higher-order categories includes specific values, such as self-direction, power, tradition, and universalism, resulting in a total of 19 values in the revised model [schwartz_refining_2012]. Although Schwartz’s framework is valuable, it may not fully capture the context-specific values that arise in HRI, where interactions between humans and robots introduce unique ethical considerations.
graham2013moral propose a model of moral foundations, which are psychological systems underlying human moral reasoning. These foundations are divided into six categories: care/harm, fairness/cheating, loyalty/betrayal, authority/subversion, sanctity/degradation, and liberty/oppression. While this model provides insights into moral reasoning, its high-level abstractions limit its direct application in the design of technological systems, including those in HRI. As such, identifying values specifically relevant to HRI, which may not be fully encompassed by these general models, remains crucial.
Another definition of values is the one presented by scanlon2000we in the buck-passing theory. In his discourse about values, Scanlon states that “to call something good is to claim that it has other properties […] which provide such reasons” for behaving in certain ways with respect to it. The value is thus not an intrinsic property of an object or situation, but entirely depends on “such reasons”. We argue that since this definition imposes fewer constraints, it is better suited to model the nuanced dynamics taking place in HRI. Consequently, the definition of value that we adopt does not follow a universal maxim and instead is more attuned to contractualist theories [scanlon1982contractualism], which emphasise respect and consideration of individual preferences in ethical decision-making.
@@ 2.2 Stakeholder perspectives
Several studies have explored the perspectives of stakeholders in high-stakes environments, providing valuable insights into the ethical challenges of HRI. cheon2016integrating highlight the importance of including roboticists’ perspectives in their design, particularly when addressing ethical concerns. In support of this, toolkits and frameworks have been developed to assist researchers in identifying and integrating ethical values into their design processes [sadek_value-sensitive_2023]. Design methodologies, such as Design Justice [costanza2020design], seek to address biases in technology design and ensure that systems are developed in a manner that is fair and just.
@@
In the field of HCI, more concrete ethical concerns have been identified. For example, makasi_chatbot-mediated_2020 examined chatbots providing public services and identified 14 key public service values. Although some of these values are specific to the public service context, many – such as adaptability, fairness, accountability, and privacy – are more broadly applicable to AI systems, including AI-powered robots. These ethical considerations offer important insights for the design of systems that interact directly with humans, such as social robots.
In the context of robotics, fosch2020gathering analysed discussions with experts, revealing key ethical, legal, and social challenges in the deployment of social robots. The authors identified five primary areas of concern: security and privacy, legal uncertainty, autonomy and agency, employment, and the replacement of human interaction. These concerns are particularly pertinent to robots designed for social good, such as those used in healthcare, education, and domestic settings. In our work, we aim to build on these insights by identifying concerns and values specific to HRI, providing a comprehensive list of values that can guide the design of robotic systems that align with human values and societal norms.
ostrowski2022ethics conducted a comprehensive review of HRI literature over the past two decades, examining various interpretations of ethics within the field. One particularly relevant aspect of their work is the exploration of ethics in human-robot relationships. In this context, they successfully identified several key ethical themes, which we aim to further validate and expand upon in our own research.
@@ 3 Phase 1: Scoping Review
The main objective of this phase is to understand the key discussions revolving around values that have emerged in the HRI field in the last two decades (RQ1). To achieve this, we conducted a scoping review focused on identifying and extracting the concerns and values from HRI studies published since 2006, when the ACM/IEEE International Conference on Human-Robot Interaction was established [10.1145/1121241].
@@ 3.1 PRISMA process
We followed the PRISMA framework [page2021prisma], the state-of-the-art workflow for systematic literature reviews, adapted for a scoping review [tricco2018prisma]. The process includes the following steps, as shown in Figure 2: (i) identification, searching specific databases within a defined time range, (ii) screening, filtering out records whose titles and abstracts do not meet the eligibility criteria, (iii) selection, reading the full texts of the publications and excluding those that do not meet the eligibility criteria, and (iv) inclusion, selecting the final list of publications and extracting their data.
@@ 3.1.1 Search query
Based on the research question RQ1 and the scope of this work, we constructed a search query as in [catania2023conversational, spitale2023systematic]. The query is satisfied for a record if the combination of the record’s title, abstract and keywords satisfies the following: contains the term social; contains the term robot or its variations; contains any of the terms value, ethic, moral, deont- (for deontology); and contains any of the terms school, teach, education, assist, healthcare, patient, home, house, domestic.
@@ 3.1.2 Exclusion criteria
  • no physical robot was present (e.g., virtual agents are excluded);
@@ 3.1.3 Selection process
Subsequently, the retrieved publications were again split randomly into two sets and assigned to the same reviewers, who read the full texts for the final selection and data extraction. Again, if a reviewer was uncertain about including a paper, this was discussed with the other reviewer for a final decision. The final selection comprises 43 publications and a list of the publications is available in Tables 3 and 4 and also in a publicly accessible table stored in our web tool [^2] https://hri-value-compass.github.io/.
@@ 3.1.4 Data extraction and analysis
Variable Type Example Publication’s year Numerical 2010, … Application topic Categorical healthcare, education, … Authors’ affiliation country Categorical USA, Germany, Europe, … Empirical study Boolean 0, 1 Population’s country Categorical United Kingdom, … Population’s age Categorical Children, Elderly, … Transcripts Qualitative “the issues of infantilisation of elderly […]”
@@
For extracting the qualitative variables, we adopted the thematic analysis approach [braun2012thematic], which consists of the following six steps. During the data extraction phase described above, for each paper assigned to them, each of the reviewers (Step 1) familiarised themselves with the data by copying verbatim parts of the papers (transcripts) that were relevant to the scope. Subsequently, they (Step 2) generated the initial codes and (Step 3) searched for themes by collating codes. Then, together, they (Step 4) reviewed the themes, comparing their results and discussing the differences, (Step 5) named the themes into meaningful values according to the literature, and (Step 6) produced a report.
@@ 3.2 Results: value-related HRI papers
Table 3: List of papers included in our survey and corresponding values and topics extracted. Note that this table has been updated with the final values defined at the end of the design process to avoid any confusion. [Continued in Table 4].
@@
Authors Year Application Study Topics Values sharkey2010ethical 2010 Healthcare No Control, Intimacy, Social enabler, Social substitute, Tricking Agency, Connectedness, Dignity, Privacy calo2011ethical 2011 Healthcare No Attachment, Expectations, Manipulation, Meta, Social enabler, Social substitute, Tricking Autonomy, Connectedness, Dignity sharkey2012granny 2012 Healthcare No Independence, Intimacy, Safety, Social substitute, Tricking Agency, Connectedness, Dignity, Privacy, Welfare amirabdollahian2013accompany 2013 Healthcare Yes Independence, Safety, Social enabler Agency, Connectedness, Welfare amirabdollahian2013assistive 2013 Healthcare No Meta, Profiling, Safety Privacy, Welfare misselhorn2013ethical 2013 Healthcare No Degradation, Social substitute, Tricking, Wellbeing Connectedness, Dignity, Welfare van2013method 2013 Ethics No Accountability, Control, Meta Agency, Equity fridin2014kindergarten 2014 Education Yes Attachment, Profiling, Safety Autonomy, Privacy, Welfare huber2016ethical 2016 Attachment No Dependence, Identity disclosure, Profiling Agency, Dignity, Privacy ienca2016social 2016 Healthcare No Meta, Safety, Transparency Privacy, Welfare sharkey2016should 2016 Healthcare No Attachment, Identity disclosure, Profiling, Social substitute, Transparency Autonomy, Connectedness, Dignity, Privacy aldinhas2017robots 2017 Healthcare No Intimacy, Profiling, Social substitute Connectedness, Privacy draper2017ethical 2017 Healthcare Yes Access, Accountability, Control, Dependence, Expectations, Manipulation, Meta, Social substitute, Wellbeing Agency, Autonomy, Connectedness, Equity, Privacy, Welfare battistuzzi2018embedding 2018 Healthcare No Companionship, Independence, Intimacy, Tricking Agency, Connectedness, Dignity, Privacy cresswell2018health 2018 Healthcare Yes Control, Safety, Tricking Agency, Dignity, Welfare garcia2018inclusion 2018 Healthcare Yes Companionship Connectedness poulsen2018care 2018 Trust Yes Identity disclosure, Transparency Dignity, Privacy vcaic2018service 2018 Healthcare Yes Assistance, Companionship, Dependence, Profiling, Safety, Social enabler, Social substitute Agency, Connectedness, Privacy, Welfare fiske2019your 2019 Mental No Accountability, Attachment, Expectations, Manipulation, Meta, Profiling, Social substitute, Tricking Autonomy, Connectedness, Dignity, Equity, Privacy li2019perceptions 2019 Home Yes Dilemma, Meta N/A malle2020general 2020 Ethics No Meta N/A poulsen2020cybersecurity 2020 Healthcare Yes Profiling, Safety Privacy, Welfare smakman2020acceptable 2020 Education Yes Accountability, Assistance, Attachment, Control, Education, Manipulation, Mentor, Nonjudgmental, Profiling, Social substitute, Tricking Agency, Autonomy, Connectedness, Dignity, Equity, Privacy, Virtue, Welfare tuuli2020use 2020 Healthcare Yes Expectations, Meta N/A van2020teachers 2020 Education Yes Accountability, Consent, Manipulation, Social substitute, Transparency, Wellbeing Autonomy, Connectedness, Equity, Privacy, Welfare vandemeulebroucke2020ethics 2020 Ethics Yes Accountability, Bias, Control, Degradation, Dependence, Exclusion, Independence, Intimacy, Meta, Profiling, Safety, Social enabler, Social substitute, Tricking, Wellbeing Agency, Connectedness, Dignity, Equity, Privacy, Welfare vanderelst2020can 2020 Healthcare Yes Accountability, Control, Expectations, Profiling Agency, Equity, Privacy cappuccio2021can 2021 Ethics No Bias, Corruptor, Dependence, Exclusion, Expectations, Fairness, Mentor Agency, Equity, Virtue
@@
Table 4: [Continued from Table 3] List of papers included in our survey and corresponding values and topics extracted. Note that this table has been updated with the final values defined at the end of the design process to avoid any confusion.
@@
Authors Year Application Study Topics Values malinverni2021educational 2021 Education Yes Wellbeing Welfare prescott2021friends 2021 Relationship No Companionship, Degradation, Expectations, Safety, Social enabler, Social substitute, Wellbeing Connectedness, Dignity, Welfare smakman2021moral 2021 Education Yes Access, Accountability, Attachment, Consent, Education, Exclusion, Expectations, Fairness, Independence, Manipulation, Mentor, Profiling, Social substitute, Transparency, Wellbeing Agency, Autonomy, Connectedness, Equity, Privacy, Virtue, Welfare wagman2021beyond 2021 Feminism No Bias, Exclusion, Expectations, Fairness, Manipulation Autonomy, Equity baisch2023elders 2022 Healthcare Yes Attachment, Bias, Expectations Autonomy, Equity carros2022ethical 2022 Ethics Yes Dilemma, Expectations, Meta N/A felber2022concept 2022 Healthcare No Accountability, Degradation, Exclusion, Profiling, Social substitute Connectedness, Dignity, Equity, Privacy hungtechnological 2022 Healthcare Yes Consent, Social enabler Connectedness, Privacy ostrowski2022mixed 2022 Healthcare Yes Attachment, Independence, Profiling, Transparency Agency, Autonomy, Privacy street2022older 2022 Healthcare Yes Control, Corruptor, Dependence, Identity disclosure, Independence, Profiling, Social substitute Agency, Connectedness, Dignity, Privacy, Virtue wheeler2023confucianism 2022 Healthcare No Exclusion, Transparency, Wellbeing Equity, Privacy, Welfare jung2023enjoy 2023 Manipulation Yes Attachment, Companionship, Education, Exclusion, Manipulation, Transparency Autonomy, Connectedness, Equity, Privacy, Welfare rosero2023using 2023 Trust Yes Dilemma, Identity disclosure Dignity soares2023ethics 2023 Healthcare Yes Accountability, Expectations, Independence Agency, Equity yeung2023living 2023 Ethics Yes Mentor, Nonjudgmental, Social enabler, Social substitute, Wellbeing Connectedness, Virtue, Welfare